OpenStudy (anonymous):

The radius of a sphere is doubled . By what factor does the surface area of the sphere change ? A. 1/4 B. 1/2 C. 2 D. 4

OpenStudy (anonymous):

@jazy

OpenStudy (amistre64):

might want to write the formula for surface area

OpenStudy (amistre64):

Area i believe has r^2 in it; so (2r)^2 pops out a 2^2 factor

OpenStudy (anonymous):

\[A=4\pi r^2\]

OpenStudy (anonymous):

Okay . Im still lost .

OpenStudy (anonymous):

how do I know the ratio of the volume ?

OpenStudy (anonymous):

Sorry I meant: The ratio of the Area = ratio of sides, squared. You know that the measures are doubled. So 1:2 is the original ratio. Now square it. \[\large 1^2 : 2^2\]

OpenStudy (anonymous):

2:4

OpenStudy (anonymous):

No, 1^2 = 1 * 1 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Right, so whats the ratio?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

The slant heights of two simlilar pyramids are in the ratio 1:5 . The volume of the smaller pyramid is 60 m ^3 . What is the volume in cube meters of the larger pyramid ?

OpenStudy (anonymous):

@Noemi95

OpenStudy (anonymous):

The ratio of 1:5 means that for every one large pyramid, 5 small pyramids make up the length of a dimension. So in this case, the height, width and depth of the larger pyramid are all 5 times longer than the small pyramid. So we can multiply the change in dimensions together to work out how much the volume has changed. 5 * 5 * 5 = 125. Here we see the volume of the large pyramid is 125 times larger than the smaller one. The volume of the smaller pyramid is 60, so: 60 * 125 = 7500.

OpenStudy (anonymous):

Thank you so much .

OpenStudy (anonymous):

No problem Lexi!(: